OpenStudy (anonymous):

What is an equation of the circle with center (-5,4) and a radius of 7? (1) (x-5)^2+ (y+4)^2=14 (2) (x-5)^2 +(y+4)^2=49 (3) (x+5)^2+ (y+ 4)^2=14 (4) (x+5)^2+ (y+ 4)^2=49

OpenStudy (ipwnbunnies):

General equation is (x-h)^2 + (y-k)^2 = r^2 Where (h,k) is the point where the center of the circle is, and 'r' is the radius of the circle. D might be your answer, but you wrote it wrong, as you told me. :3

OpenStudy (jdoe0001):

\(\bf (x-{\color{blue}{ h}})^2+(y-{\color{brown}{ k}})^2={\color{olive}{ r}}^2\qquad center\ ({\color{blue}{ -5}},{\color{brown}{ 4}})\qquad radius={\color{olive}{ 7}}\)

OpenStudy (anonymous):

Thanks